# GoDe: Gaussians on Demand for Progressive Level of Detail and Scalable Compression

### v1

@@ 4.1 Setup
The code was developed in Python 3.9 using PyTorch 2.1 and executed on a single NVIDIA A40 GPU. We set $L=8$ for all the configurations, $c_{\text{min}}=100,000$. The fine-tuning phase comprises $F=30,000$ iterations. All the other parameters remain unchanged from the original implementation. Our method was tested with three popular 3D Gaussian Splatting models: 3DGS [17], MCMC-3DGS [19], a variant of 3DGS that treats 3D Gaussians as samples from a Markov Chain Monte Carlo process, generally yielding slightly better results than 3DGS, and the neural-based Scaffold-GS [24]. For Scaffold-GS, due to its unique architecture, we adopted a slightly different compression scheme. This approach is similar to our proposed GoDe method but excludes SHs culling since SHs are not present in the model. Instead, it incorporates iterative masking with quantization-aware fine-tuning and entropy coding. The iterative masking process follows a similar approach; however, rather than masking Gaussians, we mask anchor points directly, evaluating the gradients of their respective features. Quantization is performed during the fine-tuning phase: features are quantized to 8-bit unsigned integers, while the scaling, offset, and positional attributes are quantized to half-precision floating points similarly to [30] After fine-tuning, the model parameters are further compressed using ZIP, which combines LZ77 compression with Huffman entropy coding to achieve efficient and lossless storage reduction.
Table 1: Comparison of our methods against recent methods. All methods have been retrained and tested on the same hardware. For clarity, we present four compression levels out of 8 for our configurations, selected to match the compression rate of the other methods. These levels, labeled as Ours, are derived from the same model without the need for retraining (in contrast to the other approaches). Red, orange and yellow respectively indicates the first, second and third best result.
Method Mip-NeRF360 Tanks&Temples Deep Blending
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size[MB] $\downarrow$ W $\downarrow$ FPS $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size[MB] $\downarrow$ W $\downarrow$ FPS $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size[MB] $\downarrow$ W $\downarrow$ FPS $\uparrow$
HAC (low)  [6] 27.78 0.812 0.277 22.0 502 68 24.37 0.853 0.215 11.1 307 88 30.34 0.909 0.329 6.3 190 166
HAC (high)  [6] 27.53 0.807 0.290 15.2 426 72 24.05 0.846 0.222 7.9 300 85 29.98 0.904 0.340 4.1 188 173
@@
Comp-GS [22] 27.28 0.802 0.283 16.4 485 165 23.62 0.837 0.248 9.8 241 286 29.89 0.904 0.336 10.0 246 253
27.75 0.810 0.284 19.7 453 98 24.01 0.846 0.226 11.0 266 142 30.34 0.909 0.331 8.4 63 212
27.59 0.803 0.297 14.1 369 107 23.95 0.843 0.234 8.9 230 156 30.29 0.908 0.334 7.2 56 208
27.34 0.793 0.316 10.3 288 127 23.85 0.837 0.244 7.3 196 176 30.14 0.905 0.343 5.3 44 225
Ours w/ Scaffold-GS 27.07 0.780 0.336 7.5 221 146 23.72 0.830 0.258 5.9 165 189 29.76 0.897 0.358 3.9 33 236
27.42 0.815 0.263 43.9 1546 146 23.97 0.842 0.220 22.2 938 216 29.71 0.901 0.323 19.5 931 263
27.23 0.804 0.289 20.0 603 222 23.76 0.831 0.241 11.8 438 311 29.70 0.901 0.326 11.0 491 345
26.99 0.790 0.312 13.4 381 268 23.49 0.821 0.259 8.6 301 366 29.66 0.901 0.331 8.3 357 391
Ours w/ 3DGS-MCMC 26.59 0.772 0.339 9.0 242 317 23.06 0.806 0.281 6.3 208 426 29.56 0.897 0.348 4.8 189 488
27.27 0.807 0.273 40.5 1546 189.1 23.76 0.839 0.231 21.1 938 269 29.73 0.904 0.327 19.8 931 302
27.16 0.801 0.295 18.3 603 270 23.66 0.832 0.245 11.6 438 383 29.73 0.903 0.334 10.1 491 425
26.93 0.791 0.315 12.0 381 317 23.48 0.824 0.259 8.6 301 439 29.74 0.902 0.340 7.2 357 492
Ours w/ 3DGS 26.55 0.773 0.341 7.9 242 371 23.13 0.812 0.279 6.4 208 509 29.67 0.899 0.348 5.1 189 564
@@ 4.2 Quantitative Results
Figure 3: Rate-distortion curves (left), PSNR/W (center, where W is the number of primitives) and PSNR/FPS (right) curve on Mip-NeRF360. Our approach achieves competitive results with both neural-based Gaussian models and explicit ones while maintaining significantly higher speed at comparable quality. It is also the only method that achieves scalability without requiring ad-hoc training.
We evaluated our method on the same datasets used by 3DGS, namely: Mip-NeRF360 [3], TanksAndTemples [20], and DeepBlending [15]. For each test, we report image quality metrics (PSNR, SSIM, LPIPS, computed on VGG), the model size in megabytes (SIZE), the number of primitives $|\mathcal{G}_{L-1}|$ (here $W$ for brevity), and the average frame rate on the test set (FPS). We present the results for each architecture, comparing them with various state-of-the-art proposals. To ensure a fair comparison, all the methods were re-trained and evaluated on our machine, for known discrepancies in reported results as noted by [4]. Additionally, some methods were re-trained using updated hyperparameters to generate rate-distortion curves in alignment with the parameters specified in their respective papers. The main results are summarized in Tab. 1. In Fig. 3, we plot rate-distortion and PSNR/FPS curves. These findings show that our method is generally competitive with current state-of-the-art approaches while benefiting from being a scalable model, as our rate/distortion curve is generated instantly. We divide our analysis of the results between neural-based 3DGS models and explicit ones, as the former generally provides superior rate-distortion performance.
@@
To the best of our knowledge, HAC [6] represents the current state-of-the-art in terms of rate distortion. HAC uses the Scaffold-GS [24] codebase. Scaffold-GS provides slightly higher reconstruction quality than 3DGS (approximately +0.3 dB) while reducing storage by about 5$\times$. By applying our method to Scaffold-GS we observe that it delivers comparable results, slightly outperforming HAC on the MipNeRF-360 dataset, while HAC remains superior on the Tanks&Temples dataset. We consider our results competitive, as our scalable compression method does not require re-training, unlike other methods that need to train a separate model with different hyperparameters. Our method offers another advantage beyond scalability: by masking a higher number of anchor points on average, our model is faster, achieving a frame rate increase ranging from +33% to +62% while maintaining similar reconstruction quality.
@@
Regarding the comparison with explicit Gaussian splatting models, our results are state-of-the-art on nearly every dataset in terms of rate/distortion and performance at equivalent or even superior quality, achieving up to a 3$\times$ increase in FPS. There is one point where Comp-GS [22] slightly outperforms our method, but this approach scales less effectively at higher compression rates. In this case, our approach also offers the advantage of utilizing a single scalable model, which eliminates the need for re-training when adapting to different compression rates. Notably, the models obtained with our method are highly scalable, ranging from an extremely high compression rate of a few megabytes to an extremely low one.
4.3 Visualizing LoDs
In Fig. 4 we visualize some LoD levels for various scenes. For each LoD level, we report the average frame rate computed for rendering the image of the test set, the number of Gaussians, the average PSNR, and the model size. The transition between LoDs is smooth thanks to the fine-tuning process with random LoD sampling. Moreover, we observe that lower LODs are extremely fast and small, with rendering speeds exceeding 800 FPS and model sizes as low as 2.4MB for the drjohnson scene from the DeepBlending dataset (in comparison with the vanilla-3DGS which renders the same scene at 146FPS and requires 780MB of storage). Similar performances are achieved across other datasets. It is important to note that lower LoDs tend to lose fine details, particularly in high-frequency areas such as reflections, transparency handling, and background elements. This is particulary noticeable in the kitchen scene, where at lower levels of detail the glass bottle in the background is either not rendered at all or only partially and becomes progressively more defined at higher levels. Another example can be seen in the counter scene, with the reflection on the dishes. Other examples are the paining in the drjohnson scene and the trees in the background of the $truck$ scene. This is a significant result, demonstrating the effectiveness of the method, which automatically learns a hierarchy and preserves the most important details at each level, minimizing the loss of detail to secondary aspects. This is further illustrated in Fig. 5, where we render only the Gaussian associated with an enhancement level. As the level increases, we observe that the Gaussians focus on more complex geometry and finer details.
@@ 4.3 Visualizing LoDs
Figure 5: Visualization of enhancement levels. We render only the Gaussians belonging to the specified enhancement level.
@@ 4.4.2 Pruning methods
Figure 6: Ablation on the number of levels on Mip-NeRF360.
Figure 7: Comparison among pruning methods on Mip-NeRF 360.
Figure 8: Comparison among our method with Scaffold-GS, a non-scalable variant of the same method and HAC. This constitutes an upper bound to our scalable method.
@@
LoD W Mask Only Fine Tuning Fine-Tuning + Compression
@@ 4.4.3 Effects of Fine-Tuning and Compression
We analyze the effects of fine-tuning and compression from both a quantitative and qualitative perspective. As shown in Tab. 2, the fine-tuning phase significantly improves reconstruction quality, particularly for lower LoDs, while slightly reducing quality in the two highest LoDs, of about 0.02 dB. Regarding compression, we observe a reduction in storage ranging from $99.37\%$ to $89.69\%$, depending on the level considered. The impact on reconstruction quality is minimal, with a loss of 0.03 dB to 0.18 dB across different levels. In Fig. 9, we also visualize the visual impact of fine-tuning. Non-fine-tuned models exhibit holes and generally lower quality, while the LoDs after fine-tuning show improved quality with way fewer artifacts.
@@ 4.4.4 Impact of hierarchical representation
In this ablation study, we evaluate our hierarchical compression method against a non-hierarchical variant to assess how much our hierarchical approach diverges from the best achievable performance. Specifically, we re-run our method but omit any hierarchy constraints. The process is as follows: we accumulate gradients for one epoch, mask the $k$ Gaussians with the lowest gradient values, perform a fine-tuning cycle of $7,000$ iterations, and evaluate the performance (each evaluation is a point in Fig. 8). This procedure is repeated across 8 levels, down to a minimum threshold of $100,000$ parameters, with each level acting as an independent model. Interestingly, our method sets a new performance benchmark, outperforming HAC and underline the effectiveness of iterative, gradient-based pruning for this task. Our hierarchical version, though also superior to HAC, performs less well, particularly at high compression levels. This is due to the imposed hierarchy, which requires optimizing each level in relation to the others, thus introducing a trade-off between higher and lower levels.
5 Limitation and Future Work
Our approach demonstrates effectiveness in scalable compression tasks, matching or even surpassing the performance of non-scalable compression methods. However, it also inherits certain limitations typical of top-down pruning and compression approaches, such as the requirement for a pre-trained model with a well-established geometric prior. In future work, we intend to explore the potential for a bottom-up approach, integrating the hierarchical structure during the training of the Gaussian model.

### v2

@@ 4.1 Setup
The code was developed in Python 3.9 using PyTorch 2.1 and executed on a single NVIDIA A40 GPU. We set $L=8$ for all the configurations, $c_{\text{min}}=100,000$ and $\lambda=10^{-2}$. The fine-tuning phase comprises $F=30,000$ iterations. All the other parameters remain unchanged from the original implementation. Our method was evaluated with four popular 3D Gaussian Splatting models: 3DGS [18], MCMC-3DGS [20], a variant of 3DGS that treats 3D Gaussians as samples from a Markov Chain Monte Carlo process, generally yielding slightly better results than 3DGS, the neural-based Scaffold-GS [25] and the neural and dynamic-LoD-based Octree-GS [34]. The spherical harmonic coefficients are quantized to 8-bit integers, while opacity, scaling, and rotation parameters are quantized to 16-bit floating-point values. Position parameters remain unquantized. For Scaffold-GS and Octree-GS the implementation is similar to the explicit models. Specifically, we quantize the anchor point features to 8-bit integers, while all the other parameters are quantized to 16-bit floating point values. To build the hierarchy, we use the gradient accumulated at each anchor point, constructing a progressive hierarchy of anchor points. After fine-tuning the model parameters are compressed using the LZMA algorithm.
Table 1: Comparison of our methods against recent methods. All methods have been retrained and tested on the same hardware. For clarity, we present three compression levels out of 8 for our configurations, selected to match the compression rate of the other methods. These levels, labeled as Ours, are derived from the same model without the need for retraining. Red and yellow respectively indicates the first and the second best result.
Mip-NeRF360 Tanks&Temples Deep Blending
[…]
27.43 0.811 0.271 49.0 1546 197 23.75 0.84 0.230 26.6 935 230 29.77 0.905 0.325 28.9 962 327
27.30 0.804 0.293 20.6 603 282 23.65 0.834 0.244 13.6 437 337 29.78 0.904 0.333 14.8 451 465
Ours 3DGS 26.63 0.776 0.341 8.6 242 391 23.12 0.813 0.278 7.0 208 470 29.70 0.900 0.348 7.4 212 622
27.55 0.817 0.261 47.3 1546 143 24.04 0.845 0.217 26.6 938 178 29.73 0.903 0.326 26.7 965 267
27.37 0.806 0.288 19.9 603 214 23.85 0.836 0.238 13.3 438 246 29.70 0.902 0.335 13.6 452 356
Ours 3DGS-MCMC 26.75 0.775 0.338 8.4 242 299 23.22 0.812 0.278 6.8 208 359 29.59 0.898 0.350 7.0 212 459
27.77 0.810 0.284 19.6 248 84 24.01 0.848 0.224 11.0 141 116 30.26 0.906 0.339 6.2 80 168
27.61 0.803 0.297 14.1 179 94 23.96 0.844 0.231 8.9 128 128 30.02 0.901 0.350 4.5 58 180
Ours Scaffold-GS 27.24 0.789 0.318 10.2 129 108 23.86 0.839 0.242 7.3 93 134 29.76 0.897 0.359 3.9 50 184
27.73 0.813 0.271 19.5 262 71 24.40 0.856 0.242 13.2 173 65 30.10 0.902 0.338 7.3 103 135
27.49 0.802 0.291 13.9 186 81 24.25 0.847 0.262 9.7 127 72 29.98 0.900 0.348 4.5 64 146
Ours Octree-GS 27.01 0.780 0.322 10.0 133 92 24.01 0.834 0.286 7.1 93 78 29.87 0.898 0.352 4.0 56 146
HAC (low) [7] 27.78 0.812 0.277 22.0 501 68 24.37 0.853 0.215 11.1 307 88 30.34 0.909 0.329 6.3 196 166
HAC (high) [7] 27.17 0.798 0.306 11.5 371 74 24.03 0.843 0.237 7.4 261 82 29.92 0.903 0.350 3.9 180 166
@@ 4.2 Quantitative Results
Figure 3: Rate/distortion (above) and PSNR/FPS (below) curves on the three main datasets. Our approach achieves competitive results with both neural-based Gaussian models and explicit ones while maintaining significantly higher speed at comparable quality. It is also the only method that achieves scalability without requiring ad-hoc training.
We evaluated our method on the same datasets used by 3DGS, namely: Mip-NeRF360 [4], TanksAndTemples [21], and DeepBlending [16]. For each test, we report image quality metrics (PSNR, SSIM, LPIPS, computed on VGG), the model size in megabytes (SIZE), the number of primitives $|\mathcal{G}_{L-1}|$ (here $W$ for brevity), and the average frame rate on the test set (FPS). We present the results for each architecture, comparing them with various state-of-the-art proposals. To ensure a fair comparison, all the methods were re-trained and evaluated on our machine, for known discrepancies in reported results as noted by [5]. Additionally, some methods were re-trained using updated hyperparameters to generate rate-distortion curves in alignment with the parameters specified in their respective papers. The main results are summarized in Tab. 1. In Fig. 3, we plot rate-distortion and PSNR/FPS curves. These findings show that our method is generally competitive with current state-of-the-art approaches while benefiting from being a scalable model, as our rate/distortion curve is generated instantly. We divide our analysis of the results between neural-based 3DGS models and explicit ones, as the former generally provides superior rate-distortion performance. Notably, the models obtained with our method are highly scalable, ranging from an extremely high compression rate of a few megabytes to an extremely low one.
@@
To the best of our knowledge, HAC [7] represents the current state-of-the-art in terms of rate distortion. HAC uses the Scaffold-GS [25] codebase. Scaffold-GS provides slightly higher reconstruction quality than 3DGS (approximately +0.3 dB) while reducing storage by about 5$\times$. By applying our method to Scaffold-GS and Octree-GS we observe that it delivers comparable results, slightly outperforming HAC on the Mip-NeRF360 and Tanks&Temples datasets, while HAC remains slightly superior only on the DeepBlending dataset. We consider our results competitive, as our scalable compression method does not require re-training, unlike other methods that need to train a separate model with different hyperparameters. Our method offers another advantage beyond scalability: by masking a higher number of anchor points on average, our model is faster, achieving a frame rate increase ranging from +33% to +62% while maintaining similar reconstruction quality.
@@
Regarding the comparison with explicit Gaussian splatting models, our results are state-of-the-art on nearly every dataset in terms of rate/distortion and performance at equivalent or even superior quality, achieving up to a 3$\times$ increase in FPS. There is one point where Comp-GS [23] slightly outperforms our method, but this approach scales less effectively at higher compression rates. In this case as-well, our approach also offers the advantage of utilizing a single scalable model, which eliminates the need for re-training when adapting to different compression rates.
4.3 Visualizing LoDs and Transitions
@@
In Fig. 4 we visualize some LoD levels for various scenes. For each LoD level, we report the average frame rate computed for rendering the image of the test set, the number of Gaussians, the average PSNR, and the model size. The transition between LoDs is smooth, due to the fine-tuning process with random LoD sampling. Moreover, we observe that lower LODs are extremely fast and small, with rendering speeds exceeding 900 FPS and model sizes as low as 3.7 MB for the drjohnson scene from the DeepBlending dataset (compared to the vanilla-3DGS which renders the same scene at 146FPS and requires 780MB of storage). Similar performance is achieved in other datasets. It is important to note that lower LoDs tend to lose fine details, particularly in high-frequency areas such as reflections, transparency handling, and background elements. This is particularly noticeable in the kitchen scene, where at lower levels of detail the glass bottle in the background is either not rendered at all or only partially and becomes progressively more defined at higher levels. Another example can be seen in the counter scene, with the reflection on the dishes. Other examples are the paining in the drjohnson scene and the trees in the background of the $truck$ scene. This is a significant result, demonstrating the effectiveness of the method, which automatically learns a hierarchy and preserves the most important details at each level, minimizing the loss of detail to secondary aspects. In Fig. 5, we illustrate a seamless transition between two adjacent levels of detail, achieved by linearly interpolating the opacity values of the nearest enhancement layer Gaussians. This approach further minimizes pop-up artifacts. Additional visualizations, including videos, are provided in the supplementary material.
Figure 5: Continuous transitions between two adjacent levels, with eight linear interpolating factors on Garden scene.
Figure 6: Ablation on the number of levels on Mip-NeRF360.
Figure 7: Comparison among pruning methods on Mip-NeRF 360.
Figure 8: Comparison among our method with Scaffold-GS, a non-scalable variant of the same method and HAC.
@@ 4.4.2 Pruning methods
Mask Only Fine Tuning Fine-Tuning + Compression
@@ 4.4.3 Effects of Fine-Tuning and Compression
We analyze the effects of fine-tuning and compression from both a quantitative and qualitative perspective. As shown in Tab. 2, the fine-tuning phase significantly improves reconstruction quality, particularly for lower LoDs, while slightly reducing quality in the two highest LoDs, of about 0.02 dB. Regarding compression, we observe a reduction in storage ranging from $99.37\%$ to $86.88\%$, depending on the level considered. The impact on reconstruction quality is negligible, with a loss of 0.05 dB in the highest LoD. In Fig. 9, we also visualize the visual impact of fine-tuning. Non-fine-tuned models exhibit holes and generally lower quality, while the LoDs after fine-tuning show improved quality with way fewer artifacts.
@@ 4.4.4 Impact of hierarchical representation
In this ablation study, we evaluate our hierarchical compression method against a non-hierarchical variant to assess how much our hierarchical approach diverges from the best achievable performance. Specifically, we re-run our method but omit any hierarchy constraints. The process is as follows: we accumulate gradients for one epoch, mask the $k$ Gaussians with the lowest gradient values, perform a fine-tuning cycle of $7,000$ iterations, and evaluate the performance (each evaluation is a point in Fig. 8). This procedure is repeated across 8 levels, down to a minimum threshold of $50,000$ parameters, with each level acting as an independent model. Interestingly, our method sets a new performance benchmark, outperforming HAC and underline the effectiveness of iterative, gradient-based pruning for this task. Our hierarchical version, though also superior to HAC, performs less well, particularly at high compression levels. This is due to the imposed hierarchy, which requires optimizing each level in relation to the others, thus introducing a trade-off between higher and lower levels.